A particle moving with an initial velocity of 5m/s is subjected to a uniform acceleration of - 2.5m/s2. Find the displacement in the next 4 sec. 

Grade:9

3 Answers

Saurabh Kumar
8 years ago
Use,
s = ut – ½ at2
Prajwal Kavad
8 years ago
use equations of motion : s=ut+1/2 at2 , just put the values , u=5,t=4and a=-2.5 then u will get s=10, therfore displacement is 10m
Akshat
6 years ago
s=ut+1/2at, since U=5: t=4: and a=-2.5, hence
s=5x4+1/2x(-2.5)(4)=> 20+1/2x(-2.5)x16 => 20+1/2x(-400) => 20+(-20) => 0 m
Therefore displacement would be 0m